# BIBO stability of port-Hamiltonian systems

### v1

@@ 3.5. BIBO stability of the diagonalised system
$\mathbf{G}(s)=\begin{bmatrix}\frac{1-2e^{\frac{s}{2}}}{1-2e^{s}}&\frac{1}{1-2e^{s}}\\ \frac{1}{1-2e^{s}}&-\frac{1+2e^{\frac{s}{2}}}{1-2e^{s}}\end{bmatrix}$ for which we cannot easily find the inverse Laplace transform and use it to check for BIBO stability.
@@
Now employing the decomposition from Proposition 3.13, we find $M=\frac{1}{2}\left[\begin{smallmatrix}-1&-1\\ -1&1\end{smallmatrix}\right]$. As $\|M\|_{\ell^{\infty}\rightarrow\ell^{\infty}}=1$ we cannot apply part (1) and as $\left(M_{|\cdot|}\right)^{k}=\frac{1}{2}\left[\begin{smallmatrix}1&1\\ 1&1\end{smallmatrix}\right]$ the sum $\sum_{k=0}^{\infty}\left(M_{|\cdot|}\right)^{k}$ diverges thus ruling out the use of part (2) of 3.22.
@@
implying by Theorem 3.22(3) that the pHS is BIBO stable.
@@ 4. Another approach to BIBO
Another approach to BIBO stability of pHS that we want to present here is based upon the $L^{1}$-well-posedness of these systems when considered on the state space $L^{1}$.
Lemma 4.1.
A pHS satisfying Assumptions 2.1 and 3.11 is $L^{1}$-well-posed if considered as a system on the state space $L^{1}$, that is for some (and hence all) $t>0$ there exists a constant $c>0$ such that for any classical solution $(u,x,y)$ we have that
@@
By Assumption 2.1 and [11, Thm. 1.5] the operator $A$ considered as defined analogously on a domain in $L^{1}\left([a,b],\mathbb{K}^{n}\right)$ generates a $C_{0}$-semigroup on $L^{1}\left([a,b],\mathbb{K}^{n}\right)$ and thus by [35, Thm. 7.1] the pHS considered on the state space $L^{1}\left([a,b],\mathbb{K}^{n}\right)$ is $L^{1}$-well-posed. ∎
Remark 4.2.
@@
    Note that while [11, Thm. 1.5] is formulated only for the case of a spatially constant $P_{0}$ the proof thereof does not make use of this property and it thus also holds in the more general case considered here.
@@
    It is essential here to consider the state space $L^{1}\left([a,b],\mathbb{K}^{n}\right)$, as on any reflexive Banach space (in particular any $L^{p}$ with $1<p<\infty$) $L^{1}$-well-posedness already implies boundedness of the input operator [23, Thm. 4.2.7].
Theorem 4.3.
If a port-Hamiltonian system satisfying Assumptions 2.1 and 3.11 is exponentially stable on the state space $\widetilde{X}=L^{1}\left([a,b],\mathbb{K}^{n}\right)$ then it is BIBO stable.
@@
By Lemma 4.1 we have that the pHS is $L^{1}$-well-posed on $\widetilde{X}$. But then by [23, Lem. 4.5.1] respectively an argument similar to [3, Lem. 3.22] in the general case of non-compatible systems we can show that the system is $C^{\infty}$-LILO stable in the sense of [20, Def. 4.1] and thus by [20, Thm. 4.2 & Cor. 4.4] also BIBO stable. ∎
@@
Furthermore, many approaches to checking for exponential stability, such as the Gearhart-Prüss theorem — which forms the basis of the study of exponential stability of pHS in [24] — cannot be applied in the $L^{1}$ case as they only hold on Hilbert spaces.
Results of a general nature implying exponential stability of a system on $L^{1}$ seem to exist only in restricted cases such as e.g. for positive semigoups [16, Thm. 3.5.3].
@@ 5.1. Some conjectures
During our studies of BIBO stability for distributed port-Hamiltonian systems underlying this work, a few observations have been made that suggest certain conjectures.
First, we have so far not encountered a system that satisfies the condition (2.2) as a strict inequality and is not BIBO stable. Considering how (2.2) relates to the norm of $M$ and the role this matrix plays in the construction of the counterexample in Proposition 2.5 as well as in the arguments in [24, Thm. 5.6], motivates us to propose the following conjecture on a sufficient condition for BIBO stability.
Conjecture 5.1.
If $\widetilde{W}_{B}R_{0}^{-1}\Sigma\left(\widetilde{W}_{B}R_{0}^{-1}\right)^{*}>0$ the pHS is BIBO stable.
Note that by Remark 3.4 this condition is — in the case that Assumption 3.5 holds — equivalent to $\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$.
That the condition from this conjecture, could only be a sufficient condition for BIBO stability follows from the following example, which is based upon a class of systems discussed in other contexts in [7, Sec. 5] and [24, Ex. 4.4].
Example 5.2.
Consider the system $\textrm{pH}_{2}\left(\mathbbm{1}_{2},0,\mathbbm{1}_{2},\widetilde{W}_{B},\widetilde{W}_{C}\right)$ shown to be BIBO stable in the proof of Proposition 2.6, but for which $\widetilde{W}_{B}R_{0}^{-1}\Sigma\left(\widetilde{W}_{B}R_{0}^{-1}\right)^{*}\ngtr 0$.
One potential way to approach Conjecture 5.1 — at least in the case of the diagonalised system and $P_{0}^{D}=0$ — may be via Theorem 3.22(3). Indeed, we have not yet found an example for a system with $\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$ for which no $k_{0}$ as required in this proposition exists. This motivates us to propose the following conjecture which would imply Conjecture 5.1 for this subclass of pHS.
Conjecture 5.3.
@@
Therefore we find $\left\|\left(M\mathbf{U}(s)\right)^{k}\right\|_{\mathcal{M}}=\frac{1}{2^{k}}\sum_{i=0}^{k-1}\binom{k-1}{i}\begin{bmatrix}1&1\\ 1&1\end{bmatrix}=\frac{1}{2}\begin{bmatrix}1&1\\ 1&1\end{bmatrix},$ so that
$\sum_{r}\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,r}\right\|_{\mathcal{M}}=1$ for all $k$.
Similar, we have so far not even yet encountered an exponentially stable pHS that is not also BIBO stable. This may motivate one to suggest the following — due to [12, Lem. 9.1.4] in some sense more general — conjecture.
Conjecture 5.5.
Any exponentially stable pHS is BIBO stable.
@@ 5.2. Outlook
While the final result of the approach discussed in Section 3 only provides conditions for a restricted case $P^{D}_{0}=0$ and thus in particular for the case of $P_{0}=0$ and a constant Hamiltonian, one could hope that these can be straightforwardly extended to the more general case using simple perturbation arguments as is e.g. done in [12] to show $L^{2}$-well-posedness for pHS or in [21] to show BIBO stability of certain parabolic systems. Unfortunately there are convincing reasons that this does not work analogously in this case, such as the observation that the simple shift-semigroup with identity input is not BIBO stable [20, Ex. 3.1].
Besides the extension of the results to the case $P^{D}_{0}\neq 0$ there are two further evident routes for possible generalisation. The first would be to consider higher order pHS such as they were for example investigated for asymptotic stability in [32] employing an approach also based upon an adapted concept of an associated fundamental solution. The second lies in the extension to pHS of higher spatial dimensions as were studied e.g. in [29, 22]. Here, however, one has to note that the input and output spaces may be infinite-dimensional — as with boundary control along a one-dimensional boundary for instance. In this case the characterisation of BIBO stability by the transfer function employed in the first approach is no longer valid. Hence in this case only the second approach using $L^{1}$-well-posedness may lead to sufficient conditions.

### v2

@@ 3.5. BIBO stability of the diagonalised system
    Consider $\Sigma_{2}\left(\mathbbm{1}_{2},0,2\mathbbm{1}_{2},\left[\begin{smallmatrix}1&0&\frac{1}{2}&\frac{1}{2}\\ 0&1&\frac{1}{2}&-\frac{1}{2}\end{smallmatrix}\right],\left[\begin{smallmatrix}0&0&1&0\\ 0&0&0&1\end{smallmatrix}\right]\right)$. Then we find $\mathbf{G}(s)=\begin{bmatrix}\frac{1-2e^{\frac{s}{2}}}{1-2e^{s}}&\frac{1}{1-2e^{s}}\\ \frac{1}{1-2e^{s}}&-\frac{1+2e^{\frac{s}{2}}}{1-2e^{s}}\end{bmatrix}$ for which we cannot easily find the inverse Laplace transform and use it to check for BIBO stability.
@@
    Now employing the decomposition from Proposition 20, we find $M=\frac{1}{2}\left[\begin{smallmatrix}-1&-1\\ -1&1\end{smallmatrix}\right]$. As $\|M\|_{\ell^{\infty}\rightarrow\ell^{\infty}}=1$ we cannot apply part (1) and as $\left(M_{|\cdot|}\right)^{k}=\frac{1}{2}\left[\begin{smallmatrix}1&1\\ 1&1\end{smallmatrix}\right]$ the sum $\sum_{k=0}^{\infty}\left(M_{|\cdot|}\right)^{k}$ diverges thus ruling out the use of part (2) of 29.
@@
    implying by Theorem 29(3) that the hBCS is BIBO stable.
@@ 4. Another approach to BIBO
Another approach to BIBO stability of hBCS that we want to present here is based upon the $L^{1}$-well-posedness of these systems when considered on the state space $L^{1}$.
Lemma 31.
A hBCS satisfying Assumptions 1 and 18 is $L^{1}$-well-posed if considered as a system on the state space $L^{1}$, that is for some (and hence all) $t>0$ there exists a constant $c>0$ such that for any classical solution $(u,x,y)$ we have that
@@
By Assumption 1(5) and [18, Thm. 1.5] the operator $A$ considered as defined analogously on a domain in $L^{1}\left([a,b],\mathbb{K}^{n}\right)$ generates a $C_{0}$-semigroup on $L^{1}\left([a,b],\mathbb{K}^{n}\right)$ and thus by [46, Thm. 7.1] the pHS considered on the state space $L^{1}\left([a,b],\mathbb{K}^{n}\right)$ is $L^{1}$-well-posed. ∎
Remark 32.
@@
    Note that while [18, Thm. 1.5] is formulated only for the case of a spatially constant $P_{0}$ the proof thereof does not make use of this property and it thus also holds in the more general case considered here.
@@
    It is essential here to consider the state space $L^{1}\left([a,b],\mathbb{K}^{n}\right)$, as on any reflexive Banach space (in particular any $L^{p}$ with $1<p<\infty$) $L^{1}$-well-posedness already implies boundedness of the input operator [33, Thm. 4.2.7].
Theorem 33.
If a hyperbolic boundary control system satisfying Assumptions 1 and 18 is exponentially stable on the state space $\widetilde{X}=L^{1}\left([a,b],\mathbb{K}^{n}\right)$ then it is BIBO stable.
@@
By Lemma 31 we have that the hBCS is $L^{1}$-well-posed on $\widetilde{X}$. But then by [33, Lem. 4.5.1] respectively an argument similar to [8, Lem. 3.22] in the general case of non-compatible systems (see also [43, Sec. 2.4.4]) we can show that the system is $C^{\infty}$-LILO stable in the sense of [29, Def. 4.1] and thus by [29, Thm. 4.2 & Cor. 4.4] also BIBO stable. ∎
@@
Furthermore, many approaches to checking for exponential stability, such as the Gearhart-Prüss theorem — which forms the basis of the study of exponential stability of the systems in [34] — cannot be applied in the $L^{1}$ case as they only hold on Hilbert spaces.
Results of a general nature implying exponential stability of a system on $L^{1}$ seem to exist only in restricted cases such as e.g. for positive semigoups [25, Thm. 3.5.3].
@@ 5.1. Some conjectures
During our studies of BIBO stability for the 1-D hyperbolic boundary control systems underlying this work, a few observations have been made that suggest certain conjectures.
First, we have so far not encountered a system that satisfies the condition (2) as a strict inequality and is not BIBO stable. Considering how (2) relates to the norm of $M$ and the role this matrix plays in the construction of the counterexample in Proposition 5 as well as in the arguments in [34, Thm. 5.6], motivates us to propose the following conjecture on a sufficient condition for BIBO stability.
Conjecture 34.
If $\widetilde{W}_{B}R_{0}^{-1}\Sigma\left(\widetilde{W}_{B}R_{0}^{-1}\right)^{*}>0$ then the hBCS is BIBO stable.
Note that by Remark 11 this condition is — in the case that Assumption 12 holds — equivalent to $\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$.
That the condition from this conjecture, could only be a sufficient condition for BIBO stability follows from the following example, which is based upon a class of systems discussed in other contexts in [13, Sec. 5] and [34, Ex. 4.4].
Example 35.
Consider the system $\Sigma_{2}(\mathbbm{1}_{2},0,\mathbbm{1}_{2},\widetilde{W}_{B},\widetilde{W}_{C})$ shown to be BIBO stable in the proof of Proposition 7, but for which $\widetilde{W}_{B}R_{0}^{-1}\Sigma\left(\widetilde{W}_{B}R_{0}^{-1}\right)^{*}\ngtr 0$.
One potential way to approach Conjecture 34 — at least in the case of the diagonalised system and $P_{0}^{D}=0$ — may be via Theorem 29(3). Indeed, we have not yet found an example for a system with $\|M\|_{\ell^{2}\rightarrow\ell^{2}}<1$ for which no $k_{0}$ as required in this proposition exists. This motivates us to propose the following conjecture which would imply Conjecture 34 for this subclass of hBCS.
Conjecture 36.
@@
Therefore we find $\left\|\left(M\mathbf{U}(s)\right)^{k}\right\|_{\mathcal{M}}=\frac{1}{2^{k}}\sum_{i=0}^{k-1}\binom{k-1}{i}\begin{bmatrix}1&1\\ 1&1\end{bmatrix}=\frac{1}{2}\begin{bmatrix}1&1\\ 1&1\end{bmatrix}$, so that $\sum_{r}\left\|\left(\left(M\mathbf{U}(s)\right)^{k}\right)_{p,r}\right\|_{\mathcal{M}}=1$ for all $k$.
Similar, we have so far not even yet encountered an exponentially stable hBCS that is not also BIBO stable. This may motivate one to suggest the following — due to [21, Lem. 9.1.4] in some sense more general — conjecture.
Conjecture 38.
Any exponentially stable hBCS satisfying Assumption 1 is BIBO stable.
@@ 5.2. Outlook
While the final result of the approach discussed in Section 3 only provides conditions for a restricted case $P^{D}_{0}=0$ and thus in particular for the case of $P_{0}=0$ and a constant Hamiltonian, one could hope that these can be straightforwardly extended to the more general case using simple perturbation arguments as is e.g. done in [21] to show $L^{2}$-well-posedness for pHS or in [30] to show BIBO stability of certain parabolic systems. Unfortunately there are convincing reasons that this does not work analogously in this case, such as the observation that the simple shift-semigroup with identity input is not BIBO stable [29, Ex. 3.1].
Besides the extension of the results to the case $P^{D}_{0}\neq 0$ there are two further evident routes for possible generalisation. The first would be to consider higher order hBCS such as they were for example investigated for asymptotic stability in [42] employing an approach also based upon an adapted concept of an associated fundamental solution. The second lies in the extension to higher spatial dimensions as was studied e.g. in [39, 32]. Here, however, one has to note that the input and output spaces may be infinite-dimensional — as with boundary control along a one-dimensional boundary for instance. In this case the characterisation of BIBO stability by the transfer function employed in the first approach is no longer valid. Hence in this case only the second approach using $L^{1}$-well-posedness may lead to sufficient conditions.